FULL EXPLANATION PLS

FULL EXPLANATION PLS
Answer #1

Answer U Guo, Xe) = -60-4)2- (A2-2)2 Pristo AP 2 r =M Congrange = u(xo, xa) -2 (P.x, +P2X2 CM) > -(-4)2-(12-2) 2 (Pixo +PextePixit P2 Xq =M Prixh + =y) +2] =m Peixen per day - 40 x 2 = 0 [et fes] m + ep 282 2. = + upe - 22 Pirt Per x = MP, + UP2² - 2